# Review of Degenerate Higher Order Scalar Tensor Theories in Cosmology

### v1

@@ 4 Astrophysical bodies and the Vainshtein mechanism
Significant work has been done on the effect of DHOST theories on astrophysical objects, in particular gravitational screening, neutron stars and black holes.
@@ 4.1 The Vainshtein mechanism
Dark energy and modified gravity models can change the spacetime evolution compared to general relativity, but these modifications should not appear at small scales, where precise laboratory and solar system tests can be performed. This can be achieved by the so-called screening mechanisms. In the case of DHOST theories, the Vainshtein screening [40] is at play. It has been shown to work well for Horndeski theories [41, 42, 43], whereas in the case of beyond-Horndeski theories the standard gravity is modified inside matter [44]. This new behaviour has been studied for various astrophysical objects, as presented in [45]. For DHOST theories, this topic is discussed in general in [31, 46, 47], with specific examples in [48, 49, 50, 51].
@@
Outside the source, $m^{\prime}=m^{\prime\prime}=0$ and hence the standard Newtonian potential is recovered $\Phi=\Psi=G_{\rm N}m/r$. Inside the source, the behaviour is modified, similarly to the beyond-Horndeski scenario [44].
The screening in DHOST theories has been studied in the context of the LIGO/Virgo observations in [52], and stringent constraints on the $\beta_{1}$ parameter are placed. In [48] a mechanism on evading these gravitational wave constraints is proposed.
@@ 4.2 Newtonian and Neutron stars
These results can be used to investigate the impact of such theories on various astrophysical bodies. In the case of Newtonian stars, there is a theoretical upper bound for function $\Xi_{1}$ [53].
$\Xi_{i}>-\frac{1}{6}$ (4.9)
@@
Here the matter appearing the Lagrangian $L_{m}$ is modelled by a perfect fluid with energy density $\rho$, pressure $P$ and 4-velocity $u^{\mu}$. The energy-momentum tensor is thus
@@
The deviations from general relativity $\Xi_{1}$ and $\Xi_{2}$ are given by [31]
@@
Outside the star $X=-q^{2}$ and hence the modification is indistinguishable from general relativity. Inside the star, their radial profiles must be considered, which depend on their central density and on the equation of state. This can be achieved by numerically integrating the equations of motion. In [58] several scenarios have been considered and it is shown that in DHOST theories neutron stars can have masses and radii significantly larger than in general relativity, which could provide explanations for several observations, such as the pulsar PSR J1614-2230 [60] or the compact object measured by GW190814 event [61], which have these properties.
@@
Outside the star, the results are the same as in general relativity. Inside the star, the equation can be solved numerically for each $l$ and related to the angular momentum.
@@ 4.3 Black Holes
A renewed interest in black holes can be attributed to the LIGO/Virgo direct detection of gravitational waves from binary black hole and neutron star mergers. Black holes can also be used as probes to investigate deviations from general relativity. Hence, one has to make sure that theories of modified gravity or dark energy yield constraints compatible with black holes. In standard general relativity, the most general axisymmetric stationary rotating black hole solution is the Kerr solution [64, 65]. Indeed, in the binary black hole observation, it is thought that the merger involved two spinning Kerr black holes [66]. These Kerr black holes can be generalised for DHOST theories and solutions that have non-trivial scalar field (stealth hair) have been derived [67]; linear perturbations are discussed in [68]. Generating methods for solutions by using disformal transformations are discussed in [69, 70], or by using perturbations about the flat Minkowski metric $\eta_{\mu\nu}$ along a null direction (the Kerr-Schild ansatz) in [71, 72].
We present here the Kerr solutions from DHOST theories, and we mention at the end of the section other recent developments in the field. Quadratic shift-symmetric DHOST theories with $c_{T}=1$, where all the functions in the Lagrangian only depend on $X$ are considered, such that $X=X_{0}$ is a constant. Moreover, stealth solutions are sought (solutions where the metric $g_{\mu\nu}$ is also a solution of the vacuum Einstein equations with a cosmological constant). Although such solutions are plagued by the strong coupling problem [73, 74], a controlled detuning, the scordatura mechanism [75], can be employed to solve the issue. The assumptions imply that
@@
The metric that is obtained is the usual Kerr solution in general relativity, which can be expressed in Boyer-Lindquist coordinates $(t,r,\theta,\psi)$, as
@@
where $A$ and $B$ are the disformal potentials, depending on $X$. DHOST theories are stable under these transformations [13]. Hence, assuming that the tilded action satisfies the stealth conditions (4.28), with a Kerr solution (4.32), the conditions for the untilded conditions become
@@
where the disformal transformation is constant ($A(X)=A(X_{0})\equiv A_{0}$, $B(X)=B(X_{0})\equiv B_{0}$) and the derivatives are evaluated at $X_{0}$. Imposing (4.3), these become
@@
which leads to a new Kerr solution. To avoid pathological behaviour at $r\to\infty$ (non-asymptotically flat metric at infinity), a vanishing cosmological constant is required, implying $\eta=1$, $\Theta=0$ and $E=m$ as before.
The scalar field solution becomes
@@
Hence, the disformal transformation of the stealth Kerr solution considered yields a non-stealth exact solution which is parametrised by an additional deformation parameter $\alpha$ which encodes precisely the deviations from general relativity, in addition to the mass and angular momentum parameters $(M,a)$ of the Kerr family. The sign ambiguity in (4.3) can be absorbed by redefining $t\to\pm t$ and $a\to\pm a$.
@@
This new solution is not circular [76], but similar to other Kerr black holes, it does not have a physical singularity.
In [77] it is shown that LISA can be in principle used to distinguish the Kerr geometry in DHOST from general relativity by measuring the deformation parameter $\alpha$ which quantifies departures from the standard Kerr black hole spacetime through an accumulated effect. This parameter can also be investigated by looking at its effect on the black hole shadow [78] using the observations from the Event Horizon Telescope [79, 80].
A different class of black holes solutions is the non-rotating Schwarzschild one, which can be obtained from quadratic DHOST theories [81, 74, 82], but some are disfavoured through a perturbation analysis. Perturbations to such black holes are investigated in [83, 84, 85, 86]. The strong coupling of the solutions in the DHOST theories is discussed [75], together with a solution – adding small corrections (scordatura) to avoid this issue and the implication for the black hole solutions [87]. Various generalisations have been considered, e.g. non-Schwarzschild solutions with a non-constant kinetic term [88].
Black hole solutions have been studied [89] for shift-symmetric DHOST theories which admit exact black hole solutions with a linearly time-dependent scalar field. In this case, only purely quadratic theories are viable and exact static and spherically symmetric (Schwarzschild and Schwarzschild-(anti-)de Sitter) are determined.
@@
4.4 Other astrophysical constraints
In this section we briefly introduce some of the additional effects of DHOST theories from the late universe that have been studied and that can lead to observable effects
  • The modification of the sound speed in the atmosphere of the Earth, due to the significant modification of the propagation of matter perturbations.[93]. This was used to constrain the parameters of the DHOST model.
  • Constraints on DHOST models from weak lensing, X-ray and Sunyaev - Zel’dovich observables [94, 95].
  • Constraints from large-scale structure, using two and three-point statistics with the BOSS data [96].
  • The use of helioseismology to place constraints on the parameters of the DHOST models, using the Sun as a laboratory [97, 98].
  • The effect of DHOST on structure formation through modifications in the matter power spectrum [99] and the increased sensitivity on the short-wavelength behaviour of the linear power spectrum.
@@ 5 Inflation
The inflationary paradigm, initially designed solve several cosmological paradoxes, including the horizon problem, the flatness problem and the magnetic monopole problem, is also responsible for the formation of the structures through the growth of primordial quantum fluctuations. Cosmic Microwave background (CMB) experiments, such as COBE [100], WMAP [101] and Planck [102] have tightened constraints on the observable effects on inflation, for which the main constraints are based on measurements of the angular power spectrum of CMB anisotropies and on adiabatic scalar perturbations. This exhibits a small departure from scale invariance that can be quantified at the power spectrum level through
@@
where $k_{*}$ is a pivot scale and $n_{s}$ is the scalar spectral index. Further constraints can be determined by expanding this further as
@@
$-11.1<f_{\rm NL}^{\rm local}$ $<9.3$ (5.10)
$-120<f_{\rm NL}^{\rm equil}$ $<68$ (5.11)
$-86<f_{\rm NL}^{\rm orth}$ $<10$ (5.12)
@@
To resolve this issue, one can consider Eq. (5.13) as the background and an axion-like potential [110] as a perturbation
@@
The determination of the power spectrum has been performed in [106], based on [75], following the standard formalism of inflationary perturbation theory [111]. Here we illustrate the procedure for the Lagrangian without the perturbations $m^{2}$ and $\lambda$, with the full details given in [106]. The action is perturbed at first order in perturbations and expanded up to quadratic order, and then after suitable integrations by parts, the second order DHOST Lagrangian is obtained in terms of the comoving curvature perturbation $\zeta$,
@@
During the inflationary epoch, as solutions linear in time are sought, $X=0$ and hence the $\alpha_{i}$ and $\beta_{i}$ functions become constant. There are five constant functions (and an additional two perturbation parameters) that can be chosen to match the CMB observations.
In the case of the bispectrum [107], one can employ the formalism described in [115, 116]: expanding the action at third order in the comoving gauge, computing the Hamiltonian in the interaction picture and employing the in-in formalism to determine the three-point correlation function. In addition to the parameters fit for the power spectrum, there are another five that appear in the bispectrum.
For the power spectrum, parameters that match the CMB observations can be readily obtained, but for the bispectrum this is more involved due to its three-dimensional nature, requiring a full CMB likelihood analysis to confront to the Planck data. Parameters compatible with both the powers spectrum and the bispectrum have been found in [108].
6 Alternatives to inflation
Another idea related to the use of DHOST models in the early universe is presented in [117, 118, 119, 120], where an alternative to inflation – the bouncing cosmology [121, 122, 123, 124] – is built using these theories.
This model is inspired by the Galilean Genesis model [125, 126] and aims to resolve several potential problems such as no-go theorems, strong coupling issues and connections to standard cosmology.

### v3

@@ 4 Astrophysical Bodies and the Vainshtein Mechanism
Significant work has been performed on studying the effects of DHOST theories on astrophysical objects, in particular gravitational screening, neutron stars and black holes.
@@ 4.1 The Vainshtein Mechanism
Dark energy and modified gravity models can change the spacetime evolution compared to GR, but these modifications should not appear at small scales, where precise laboratory and solar system tests can be performed. This can be achieved by the so-called screening mechanisms. In the case of DHOST theories, the Vainshtein screening [74] is at play. It has been shown to work well for Horndeski theories [75, 76, 77], whereas in the case of beyond-Horndeski theories the standard gravity is modified inside matter [78]. This new behaviour has been studied for various astrophysical objects, as presented in [79]. For DHOST theories, this topic is discussed in general in [62, 80, 81], with specific examples in [82, 83, 84, 85].
@@
The parameter $\alpha_{V}$ quantifies the non-linear action [86] and is defined in terms of the functions appearing in the action as
$\alpha_{V}={4X(F_{2,X}-2A_{2}-2XA_{2,X})}/{M^{2}}\,.$ (4.9)
@@
Outside the source, $m^{\prime}=m^{\prime\prime}=0$ and hence the standard Newtonian potential is recovered, $\Phi=\Psi=G_{\rm N}m/r$. Inside the source, the behaviour is modified, similarly to the beyond-Horndeski scenario [78].
The screening in DHOST theories has been studied in the context of the LIGO/Virgo observations in [87], and stringent constraints on the $\beta_{1}$ parameter are placed. Mechanisms to evade these gravitational waves constraints have been proposed in [88, 67, 82].
@@ 4.2 Newtonian and Neutron Stars
These results can be used to investigate the impact of such theories on various astrophysical bodies. In the case of Newtonian stars, there is a theoretical lower bound for the function $\Xi_{1}$ [89],
$\Xi_{1}>-\frac{1}{6}\,.$ (4.10)
@@
Here the matter appearing in the Lagrangian $L_{m}$ is modelled by a perfect fluid with energy density $\rho$, pressure $P$ and 4-velocity $u^{\mu}$. The energy-momentum tensor is thus
@@
The deviations from GR $\Xi_{1}$ and $\Xi_{2}$ are given by [62]
@@
Outside the star $X=-q^{2}$ and hence the modification is indistinguishable from GR. Inside the star, their radial profiles must be considered, which depend on their central density and on the equation of state. This can be achieved by numerically integrating the equations of motion. In [94] several scenarios have been considered and it was shown that in DHOST theories neutron stars can have masses and radii significantly larger than in GR, of up to 50% depending on the numerical values considered for the parameters, which could provide explanations for several observations, such as the pulsar PSR J1614-2230 [96] or the compact object measured by GW190814 event [97], which have these properties.
@@
with $w_{l}$ satisfying
$\frac{(hf)^{\frac{1}{2}}}{r^{4}F}\frac{d}{dr}\left[\frac{r^{4}F}{(hf)^{\frac{1}{2}}}w_{l}^{\prime}\right]+\frac{h((l+1)l-2)}{r^{2}}w_{l}-\frac{h(P+\rho)}{F}(\Omega+w_{l})=0\,.$ (4.29)
Outside the star, the results are the same as in GR. Inside the star, the equation can be solved numerically for each $l$. For $l=1$ one gets,
$w_{1}^{\prime\prime}+\frac{4}{r}w_{1}^{\prime}=0\,,$ (4.30)
which has a solution of the form
$w_{1}=-2\frac{J}{r^{3}}\,.$ (4.31)
The integration constant $J$ physically corresponds to the angular momentum. The moment of inertia is defined as $I=J/\Omega$. Then one can fix the integration constant $J$ in order to match the solutions inside and outside the neutron star. Hence, the moment of inertia of the neutron star $I$ can be expressed as
[…]
In GR, universal relations relating the normalised moment of inertia
$\tilde{I}=I/(MR^{2})$ (4.33)
and the stellar compactness
${\cal C}=\frac{M}{R}\,,$ (4.34)
as well as the dimensionless moment of inertia
$\bar{I}=I/M^{3}$ (4.35)
and the compactness ${\cal C}$ [100, 101] have been derived. These universal relations are respectively of the form:
$\frac{I}{MR^{2}}$ $=$ $a_{0}+a_{1}\,{\cal C}+a_{4}\,{\cal C}^{4}\,,$ (4.36)
$\frac{I}{M^{3}}$ $=$ $a_{-1}\,{\cal C}^{-1}+a_{-2}\,{\cal C}^{-2}+a_{-3}\,{\cal C}^{-3}+a_{-4}\,{\cal C}^{-4},$ (4.37)
where the constants $a_{i}$ can be obtained numerically.
In [94], the robustness of these universal relations has been analysed in the context of DHOST theories for various deviations from GR, showing that $\mid 1-I/I_{fit}\mid<0.1/0.05$ for relations (4.36) and (4.37) respectively.
@@ 4.3 Black Holes
A renewed interest in black holes can be attributed to the LIGO/Virgo direct detection of gravitational waves from binary black hole and neutron star mergers. Black holes can also be used as probes to investigate deviations from GR. Hence, one has to make sure that theories of modified gravity or dark energy yield constraints compatible with black holes. In standard GR, the most general axisymmetric stationary rotating black hole solution is the Kerr solution [102, 103]. Indeed, in the binary black hole observation, it is thought that the merger involved two spinning Kerr black holes [3]. These GR Kerr solutions can be generalised to DHOST theories. In [104], DHOST Kerr black hole solutions have been derived, which have stealth hair, i.e. the scalar field is non-trivial. Linear perturbations around these solutions are discussed in [105]. Methods for generating solutions by using disformal transformations are discussed in [106, 107], or by using perturbations about the flat Minkowski metric $\eta_{\mu\nu}$ along a null direction (the Kerr-Schild ansatz) in [108, 109].
We present here the Kerr solutions from DHOST theories, and we mention at the end of the section other recent developments in the field. Quadratic shift-symmetric DHOST theories with $c_{T}=1$, where all the functions in the Lagrangian only depend on $X$ are considered, such that $X=X_{0}$ is a constant. Moreover, stealth solutions are sought (solutions where the metric $g_{\mu\nu}$ is also a solution of the vacuum Einstein equations with a cosmological constant). Although such solutions are plagued by the strong coupling problem [110, 111], a controlled detuning, the scordatura mechanism [112, 113], can be employed to solve the issue. The assumptions imply that
@@
The metric that is obtained is the usual Kerr solution in GR, which can be expressed in Boyer-Lindquist coordinates $(t,r,\theta,\psi)$ as
@@
where $A$ and $B$ are the disformal potentials, depending on $X$. DHOST theories are stable under these transformations [39]. Hence, assuming that the tilded action satisfies the stealth conditions (4.38), with a Kerr solution (4.42), the conditions for the untilded equations become
@@
where the disformal transformation is constant ($A(X)=A(X_{0})\equiv A_{0}$, $B(X)=B(X_{0})\equiv B_{0}$) and the derivatives are evaluated at $X_{0}$. Imposing (4.3), these simplify to
@@
which leads to a new Kerr solution. To avoid a pathological behaviour at $r\to\infty$ (non-asymptotically flat metric at infinity), a vanishing cosmological constant is required, implying $\eta=1$, $\Theta=0$ and $E=m$ as before. The scalar field solution becomes
@@
Hence, the disformal transformation of the stealth Kerr solution considered yields a non-stealth exact solution which is parametrised by an additional deformation parameter $\alpha$ which encodes precisely the deviations from GR, in addition to the mass and angular momentum parameters $(M,a)$ of the Kerr family. The sign ambiguity in (4.3) can be absorbed by redefining $t\to\pm t$ and $a\to\pm a$.
@@
This new solution is not circular [114], but as it the case with other Kerr black holes, it does not have a physical singularity.
In [115] it was shown that LISA can be in principle used to distinguish the Kerr geometry in DHOST from GR by measuring the deformation parameter $\alpha$ which quantifies departures from the standard Kerr black hole spacetime through an accumulated effect. This parameter can also be investigated by looking at its effect on the black hole shadow [116] using the observations from the Event Horizon Telescope [117, 118].
A different class of black holes solutions is represented by the non-rotating Schwarzschild ones, which can be obtained from quadratic DHOST theories [119, 111, 120], though some are disfavoured through a perturbation analysis. Perturbations to such black holes are investigated in [121, 122, 123, 124]. The strong coupling of the solutions in the DHOST theories is discussed [112, 113], together with a solution – adding small corrections (scordatura) to avoid this issue and the implication for the black hole solutions [125]. Various generalisations have been considered, e.g. non-Schwarzschild solutions with a non-constant kinetic term [126].
Black hole solutions have been studied [127] for shift-symmetric DHOST theories which admit exact black hole solutions with a linearly time-dependent scalar field. In this case, purely quadratic theories are viable and exact static and spherically symmetric solutions (Schwarzschild and Schwarzschild-(anti-)de Sitter) are determined. It was also shown that only a subset of the cubic theories admit stealth solutions.
@@
5 Other Constraints
In this section we briefly introduce some of the additional effects of DHOST theories that have been studied and that can lead to observable effects.
Sound Speed Constraints
The modification of the Newtonian potential in DHOST theories with respect to GR induces modifications to the matter perturbations. In [131] it has been shown that these modifications can change the speed of sound in the atmosphere of the Earth and can generate instabilities in homogeneous media. It was shown that this modification can be encoded in a single parameter that can, in principle, be constrained.
Galaxy Cluster Constraints
DHOST theories modify the Newtonian potential (see Sec. 4.1) and are expected to also induce deviations from GR at galaxy clusters scales hence affecting weak lensing, X-ray and Sunyaev - Zel’dovich observables. In [132], expressions for the lensing convergence $\kappa$, and the pressure profile $P$ of clusters in the framework of DHOST theories are determined, quantifying how much they deviate from their GR counterparts. Using the galaxy cluster sample of BOXSZ [133], it has been shown that combined measurements of $\kappa$, $P$ and of the electron number density, $n_{e}$, can constrain both the cluster and DHOST parameters.
In Ref. [134], the modifications from DHOST theories were confronted to data coming from the X-COP galaxy cluster sample [135]. Using Bayesian methods, a constraint of $\Xi_{1}=-0.030\pm 0.043$ (95% confidence level) has been obtained.
Constraints from Large-Scale Structure
The large-scale structure of the universe has been used to constrain modified gravity models using two and three-point statistics [136]. In this work a joint analysis of the anisotropic components of galaxy two- and three-point correlation functions (3PCF) to actual galaxy data and has been performed, together with the computation of constraints of the nonlinear effects of DHOST theories on cosmological scales. Using the Baryon Oscillation Spectroscopic Survey (BOSS) data release 12, lower bounds of $-1.655<\xi_{t}$ and $-0.504<\xi_{s}$ at 95% confidence level on the parameters characterising the time evolution of the tidal and shift terms of the second-order velocity field were derived. These constraints are consistent with GR predictions of $\xi_{t}=15/1144$ and $\xi_{s}=0$. This represents a 35 and 20-fold improvement respectively, over using only the the isotropic 3PCF.
Helioseismology Constraints
The screening mechanism described in Sec. 4.1 provides an opportunity to perform precision tests of gravity at stellar scales. In [137, 138], the Sun is used as a laboratory where imprints of the fifth-force effect on the solar equilibrium structure are investigated. The work details how different solar regions can be used to test gravity using analytic techniques and numerical simulations. It is shown that the fifth force leaves an important signature on the solar sound speed, even when accounting for solar microphysics such as opacity, diffusion, equation of state and metallicity. The fifth-force coupling strength of has been constrained to $-10^{-3}\leq\mathcal{Y}\leq 5\times 10^{-4}$ (at 95% confidence level), improving previous results by a factor of three.
Matter Power Spectrum
In [139], matter density perturbations up to third order and the one-loop matter power spectrum in DHOST theories have been investigated. This has involved systematically solving gravitational field equations and fluid equations order by order. Three shape functions that characterise the third-order solutions in DHOST theories are presented. Convergence conditions of the loop integrals in the infrared and ultraviolet limits are discussed, showing that these constraints are more stringent than in GR. Hence, the one-loop matter power spectrum is sensitive to the short-wavelength behaviour of the linear power spectrum.
@@ 6 Inflation
The inflationary paradigm, initially designed solve several cosmological paradoxes, including the horizon problem, the flatness problem and the magnetic monopole problem, is also responsible for the formation of the structures through the growth of primordial quantum fluctuations. Cosmic Microwave background (CMB) experiments, such as COBE [140], WMAP [141] and Planck [31] have tightened the constraints on the observable effects on inflation. The tightest constraints are based on measurements of the angular power spectrum of CMB anisotropies. The measurements show that the power spectrum exhibits a small departure from scale invariance that can be quantified through
@@
where $k_{*}$ is a pivot scale and $n_{s}$ is the scalar spectral index. Further constraints can be determined by expanding this expression further as
@@
$-11.1<f_{\rm NL}^{\rm local}$ $<9.3\,,$ (6.10)
$-120<f_{\rm NL}^{\rm equil}$ $<68\,,$ (6.11)
$-86<f_{\rm NL}^{\rm orth}$ $<10\,,$ (6.12)
@@
To resolve this issue, one can consider Eq. (6.13) as the background and add an axion-like potential [149] as a perturbation
@@
The determination of the power spectrum has been performed in [145], based on [113], following the standard formalism of inflationary perturbation theory [150]. Here we illustrate the procedure for the Lagrangian without the perturbations $m^{2}$ and $\lambda$, with the full details given in [145]. It is convenient to work with dimensionless coordinates (tilded) and variables (expressed in lowercase), which can be achieved by using two mass scales, $M$ and $\Lambda\simeq m_{\rm Pl}$, defined by:
${\tilde{t}}\equiv\Lambda t\,,\hskip 28.45274pt{\tilde{x}}^{i}\equiv\Lambda x^{i}\,,$ (6.16)
[…]
The action is perturbed at first order in perturbations and expanded up to quadratic order, and then, after suitable integrations by parts, the second order DHOST Lagrangian is obtained in terms of the comoving curvature perturbation $\zeta$,
@@
During the inflationary epoch, as solutions linear in time are sought, $X=-1$ and hence the $\alpha_{i}$ and $\beta_{i}$ functions become constants. There are five constant functions (and an additional two perturbation parameters) that can be chosen to match the CMB observations. In [145], two parameter sets are found. Both sets satisfy the Planck constraints, but have different tensor-to-scalar ratios, showing that one would be able to use DHOST theories to describe inflation even if the tensor-to-scalar ratio would be significantly more accurately measured in the future with CMB experiments such as LiteBIRD [154]. These are given by
$\alpha_{B}=1,$ $\quad\alpha_{H}=1.04,\quad\beta_{K}=3.97343,\quad f_{2}=2.7,\quad h_{b}=3\times 10^{-5}$
$\lambda=10^{-36},\quad m^{2}=-1.6\times 10^{-23}\quad(r=0.04)\,;$ (6.25)
$\alpha_{B}=1,$ $\quad\alpha_{H}=1.001,\quad\beta_{K}=3.9993,\quad f_{2}=8.8,\quad h_{b}=10^{-5}\quad$
$\lambda=5\times 10^{-32},\quad m^{2}=-1.5\times 10^{-27}\quad(r=10^{-3})\,.$ (6.26)
In the case of the bispectrum [146], one can employ the formalism described in [155, 156]: expanding the action at third order in the comoving gauge, computing the Hamiltonian in the interaction picture and using the in-in formalism to determine the three-point correlation function. In addition to the parameters fit for the power spectrum, there are another five that appear in the bispectrum. Due to its 3D nature, the analysis of bispectrum is more involved. As a first order approximation, one can compare the bispectrum of the DHOST model with the three most commonly used templates from Planck – local, equilateral and orthogonal, described in Eqs. (6.7)-(6.9) in the equilateral configuration and take into account the projection of the DHOST bispectrum on these shapes [157]. This analysis shows that the additional parameters can be tuned such that the overlap between the bispectra and the three templates can be made arbitrarily small [146]. This result is, however, misleading, as a full likelihood analysis using the Planck 2018 CMB results has shown that the bispectrum consists of a fixed contribution (coming from the power spectrum parameters) and a linear combination of terms depending on five new parameters defining the cubic perturbations. They have different shapes, with the former peaking in the squeezed limit, and the latter in the equilateral limit. These shapes cannot cancel sufficiently if the amplitude of the squeezed shape is too large. It has thus been shown in [147] that the model described in Eq. (6) is not favoured by the data. By using the CMB-BEST code [158] and marginalising over the free parameters it has been shown explicitly that there are viable DHOST inflationary models satisfying both power spectrum and bispectrum constraints from Planck, with a model satisfying both power spectrum and bispectrum constraints having parameters
$\alpha_{B}=1,$ $\quad\alpha_{H}=1.04,\quad\beta_{K}=3.97343,\quad f_{2}=6,\quad h_{b}=0.00001712,$
$\lambda=1.2\times 10^{-36},\quad m^{2}=-10^{-23}\,.$ (6.27)
7 Other Developments for the Early Universe
The initial singularity problem appearing in inflation has led to the development of alternative theories. Such alternatives include bouncing cosmologies [33] and Galilean Genesis models. The instability problem in bouncing cosmology has been studied in Refs. [159, 160, 161, 162, 163, 164]. In [161], it was found for the first time that a stable bouncing cosmology can be constructed in a beyond Horndeski theory without encountering a strong coupling problem. These works have led to further investigation of alternatives to inflation within beyond Horndeski or DHOST theories.
In [165, 166, 167, 168] the DHOST Genesis model, inspired by Galilean Genesis [35, 169] is built using DHOST theories. Potential problems such as no-go theorems for bouncing cosmology [170, 171, 172, 173] have been resolved by using DHOST theories [165, 167]. Similarly, it is shown in [166, 168] that DHOST theories can also evade the no-go theorems for Galilean Genesis cosmology [35, 169].